OpenStudy (anonymous):

Simplify Square root of 224

OpenStudy (anonymous):

\[\sqrt{224}\] \[\sqrt{16*14}\] \[4\sqrt{14}\]

OpenStudy (anonymous):

\[4\sqrt{14}\]